# Equivariant scaling asymptotics for Poisson and Szegő kernels on Grauert tube boundaries

### v1

@@ Abstract
2): There are furthermore an induced action of $G$ as a group of CR and contact automorphisms on $X^{\tau}$, and a corresponding unitary representation on the Hardy space $H(X^{\tau})$. The action of $G$ on $X^{\tau}$ commutes with the homogeneous ‘geogesic flow’  and the representation on the Hardy space commutes with the elliptic self-adjoint Toeplitz operator induced by the generator of the goedesic flow. Hence each eigenspace of the latter also splits over the irreducible representations of $G$. We study the asymptotic concentration of the eigenfunctions in a given isotypical component.
@@ 1 Introduction
As proved and discussed in [GS-1991], [GS-1992], [Lem], [LS], [SZ02] (see also [PW-1989] and [PW-1991]), the choice of a $\mathcal{C}^{\varpi}$ Riemannian metric $\kappa$ on $M$ singles out a unique $\mathcal{C}^{\varpi}$ exhaustion function on $\tilde{M}$ (perhaps after passing to a smaller tubular neighbourhood of $M$ in $\tilde{M}$), with the following properties:
@@
Grauert tubes have been studied extensively in recent years, both for their intrinsic geometric interest and in relation to the following fundamental holomorphic extension property of the eigenfunctions of the non-negative Laplacian $\Delta$ of $(M,\kappa)$, which was first discovered by Bouted de Monvel [BdM-1978]; this foundational result was also proved and analyzed in [Leb], [S-2014], [S-2015],[Z-2020] (see furthermore the discussion in [GS-1992]). We give the formulation in [Leb].
Let us denote by
@@
    $\left(e^{-\tau\,\epsilon_{j}}\,\langle\mu_{j}\rangle^{(d-1)/4}\,\tilde{\varphi}_{j,k}^{\tau}\right)_{j\geq 1,\,k=1,\ldots,n_{j}}$
@@
    a privileged collection of CR functions (elements of $H(X^{\tau})$), given by the restrictions $\tilde{\varphi}^{\tau}$ to $X^{\tau}$ of the holomorphic extensions $\varphi$ of the eigenfunctions of $\Delta$ (if $\tau\in(0,\tau_{1})$);
@@
In fact, the principal symbol of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ (i.e., the restriction to $\Sigma^{\tau}$ in (6) of the principal symbol of $D^{\tau}_{\sqrt{\rho}}$) is strictly positive; hence the spectrum of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ is discrete and accumulates only at +$\infty$+ ∞ [BdM-G].
@@
The second requirement means that $0\in\mathfrak{g}^{\vee}$ is a regular value of $\left.\Phi\right|_{\tilde{M}^{\tau_{0}}\setminus M}$. It is satisfied a fortiori if $\mu$ itself is locally free. Under Assumption 4, $d>d_{G}$ (see Corollary 55 below).
Let $\Gamma^{\tau}:\mathbb{R}\times X^{\tau}\rightarrow X^{\tau}$ denote the flow of $\upsilon_{\sqrt{\rho}}$ ($\Gamma^{\tau}$ will be referred to - with some abuse of language - as the ‘homogeneous geodesic flow’,   since it is intertwined with the latter by the imaginary time exponential map - see [GS-1992], [LS], [GLS]). Since $\tilde{\mu}^{\tau}$ and $\Gamma^{\tau}$ commute, there is a product action of $G\times\mathbb{R}$ on $X^{\tau}$; our first result is that (16) and (17) asymptotically concentrate near certain compact loci $\mathfrak{X}^{\tau}_{\chi}\subset X^{\tau}\times X^{\tau}$ determined by $\chi$, the moment map, and the orbits of the latter action.
@@
We are thus led to consider the asymptotics of $P^{\tau}_{\chi,\nu,\lambda}(x,y)$ and $\Pi^{\tau}_{\chi,\nu,\lambda}(x,y)$ when $(x,y)$ ranges in a shrinking neighourhood of a fixed pair $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$. As in [P-2024], the asymptotics in point will be formulated in suitable systems of local coordinates at points $x\in X^{\tau}$, called normal Heisenberg local coordinates (in the following, NHLC’s), which are a slight specialization of the Heisenberg local coordinates in [CR1] and [CR2] (see §3.3 of [P-2024]).
@@
1.1 Scaling asymptotics in the action-free case
@@
The leading order term describing the exponential decay in the scaling asymptotics of $P^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ and $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ for $\lambda\rightarrow+\infty$ depends on a real-quadratic complex valued form on the vector subspace $\mathcal{H}_{x_{1}}^{\tau}\times\mathcal{H}_{x_{2}}^{\tau}\subseteq T_{x_{1}}X^{\tau}\times T_{x_{2}}X^{\tau}$. To describe the latter form, we need a brief digression.
@@
::=:= $\frac{1}{2}\,\left(Z_{1}^{\dagger}\,\overline{Q}\,P^{-1}\,Z_{1}+2\,\overline{Z}_{2}^{\dagger}\,P^{-1}\,Z_{1}-\overline{Z}_{2}^{\dagger}\,P^{-1}\,Q\,\overline{Z}_{2}-\|Z_{1}\|^{2}-\|Z_{1}\|^{2}\right).$
@@
If $(V,h^{\prime})$ is a Hermitian complex $k$-dimensional vector space, so that $h=g_{h}-\imath\,\omega_{h}$, where $g_{h}:=\Re(h)$ and $\omega_{h}:=\Im(h)$ are, respectively, an Euclidean product and a symplectic bilinear form on $V$. We define $\psi_{2}^{\omega_{h}}:V\times V\rightarrow\mathbb{C}$ by
@@
Let us identify the unitary group $U(d-1)$ as the maximal compact subgroup $\hat{U}(d-1):=\mathrm{Sp}(2d-2)\cap O(2d-2)$ of $\mathrm{Sp}(2d-2)$ in the standard manner. Then $A\in\hat{U}(d-1)\leqslant\mathrm{Sp}(2d-2)$ if and only if in (22) we have $Q=0$ and $P\in U(d-1)\leqslant\mathrm{GL}_{\mathbb{C}}(d-1)$. Hence if $A\in\hat{U}(d-1)$ then
$\Psi_{A}(\mathbf{v}_{1},\mathbf{v}_{2})$ $=$ $\overline{Z}_{2}^{\dagger}\,P^{-1}\,Z_{1}-\frac{1}{2}\,\|Z_{1}\|^{2}-\frac{1}{2}\,\|Z_{1}\|^{2}$
$=$ $h_{0}(Z_{1},P\,Z_{2})-\frac{1}{2}\,\|Z_{1}\|^{2}-\frac{1}{2}\,\|Z_{1}\|^{2}$
@@
$P^{\tau}_{\chi,\lambda}(x_{1,\lambda},x_{2,\lambda})$
$\sim$ $\frac{1}{\sqrt{2\,\pi}}\cdot\left(\frac{1}{2\,\pi}\right)^{d-1}\cdot\left(\frac{\lambda}{\tau}\right)^{\frac{d-1}{2}}\cdot e^{\frac{1}{\tau}\,\left[\imath\,\sqrt{\lambda}\,(\theta_{1}-\theta_{2})+\Psi_{B^{-1}}(\mathbf{v}_{1},\mathbf{v}_{2})\right]}\,e^{-\imath\,\lambda\,t_{1}}$
$\cdot e^{\imath\,\tilde{\theta}^{\tau}(x_{1},x_{2})}\cdot\left[\chi(t_{1})+\sum_{k\geq 1}\,\lambda^{-k/2}\,\widetilde{F}_{k}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2}\right)\right],$
@@
When $x_{1}=x_{2}$ and $\chi\in\mathcal{C}^{\infty}_{c}\big{(}(-\epsilon,\epsilon)\big{)}$ for suitably small $\epsilon>0$ then $M=I_{2d-2}$, $t_{1}=0$ and $\theta^{\tau}(x_{1},x_{2})=\tilde{\theta}^{\tau}(x_{1},x_{2})=0$. Thus we recover the near-diagonal scaling asymptotic in [P-2024]. Furthermore, we recover the near-graph scaling asymptotics of [CR2], with an explicit determination of the leading factor, by fixing $(\theta_{j},\mathbf{v}_{j})$ and rescaling according to Heisenberg type (that is, in the form $(\theta_{j}/\lambda,\mathbf{v}_{j}/\sqrt{\lambda})$).
1.2 Scaling asymptotics in the general equivariant case
In order to state our results in the general equivariant setting, we need to introduce a finer decomposition of $T_{x}X^{\tau}$, valid at $x\in Z^{\tau}$, than the one dictated by $\alpha^{\tau}$ in (8); this decomposition depends on $\mu^{\tau}$ and $\Phi$, and is the analogue of the decomposition used in [P-2008] in the line bundle setting.
@@ 1.2 Scaling asymptotics in the general equivariant case
Given a choice of NHLC’s centered for $X^{\tau}$ at $x$, we obtain a direct sum decomposition of $\mathbb{R}^{2\,d-1}=\mathbb{R}\oplus\mathbb{C}^{d-1}\cong T_{x}X^{\tau}$, dictated by $\alpha^{\tau}$, where $\mathbb{C}^{d-1}\cong\mathcal{H}^{\tau}_{x}$. The finer decomposition of $\mathcal{H}^{\tau}_{x}$ in Definition 18, dictated by $\tilde{\mu}^{\tau}$, determines a a corresponding decomposition of $\mathbb{C}^{d-1}$, which - emphasizing dimensions - we shall write in the form
@@ 1.2.1 Scaling asymptotics for $\Pi^{\tau}_{\chi,\nu,\lambda}$
Under Assumption 4, the effective volume at $x\in Z^{\tau}$, denoted $V_{eff}(x)$, is the volume of the $G$-orbit through $x$, $x^{G}\subset X^{\tau}$ for the induced Riemannian density (with respect to $\tilde{\kappa}$ see also Remark 14).
@@
$A_{\chi}(x_{1},x_{2})$ ::=:= $\int_{\mathbb{R}^{d_{G}}_{t}}\,\mathrm{d}\mathbf{u}^{t}\,\int_{\mathbb{R}^{d_{G}}_{v}}\,\mathrm{d}\mathbf{u}^{v}\,\int_{\mathbb{R}^{2d-2-2d_{G}}_{h}}\,\mathrm{d}\mathbf{u}^{h}$ (26)
@@
$\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})$ ::=:= $\chi(t_{1})\cdot\frac{\left|\det(P)\right|}{r\cdot V_{eff}(x_{1})}\cdot\frac{A_{\chi}(x_{1},x_{2})}{\pi^{d-1}},$
$\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$ ::=:= $\dim(\nu)\cdot\overline{\Xi_{\nu}\left(g_{l}\right)},$
@@
    if $B$ is orthogonal one also has $|\det(P)|=1$ (since in this case $P$ is a unitary complex matrix), hence $\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})=\chi(t_{1})/\big{(}r\cdot V_{eff}(x_{1})\big{)}$.
In view of Theorem 7, we expect an exponential decay of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ (and $P^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$) along normal displacements to $Z^{\tau}$, that is, when $\mathbf{v}_{j}=\mathbf{v}_{j}^{t}\in T_{x_{j}}^{t}X^{\tau}$.
@@
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,l,\nu}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right];$
@@
    $\mathbf{v}_{1}=\mathbf{v}_{1}^{t}+\mathbf{v}_{2}^{h}\in T^{t}_{x_{1}}X^{\tau}\oplus T^{h}_{x_{1}}X^{\tau},\quad\|\mathbf{v}_{1}\|\leq C\,\lambda^{\epsilon^{\prime}}$
@@
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{k\geq 1}\,\lambda^{-k/2}\,\tilde{F}_{k,l,\nu}\left(x_{1},x_{2};\theta_{1},\theta_{2},\mathbf{V}\right)\right];$
@@
The unitary factor $e^{\imath\,\theta^{\tau}_{t}(x)}$ is a smooth function on ×$\mathbb{R}\times X^{\tau}$ and is related to the description (by Zelditch) of the unitary Toeplitz operators $e^{\imath\,t\,\mathfrak{D}^{\tau}_{\sqrt{\rho}}}$ as ‘dynamical Toeplitz operators’ (recall (10)); it is therefore an intrinsic invariant, given a priori, of the CR structure of $X^{\tau}$. Similar considerations apply to the unitary factor $e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}$ appearing in the asymptotics of $P^{\tau}_{\chi,\nu,\lambda}$ of §1.2.2, with $e^{\imath\,t\,\mathfrak{D}^{\tau}_{\sqrt{\rho}}}$ replaced by $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ in (6) below - see (149) and (151) below.

### v3

@@ Abstract
2): There are furthermore an induced action of $G$ as a group of CR and contact automorphisms on $X^{\tau}$, and a corresponding unitary representation on the Hardy space $H(X^{\tau})$. The action of $G$ on $X^{\tau}$ commutes with the homogeneous ‘geodesic flow’  and the representation on the Hardy space commutes with the elliptic self-adjoint Toeplitz operator induced by the generator of the goedesic flow. Hence each eigenspace of the latter also splits over the irreducible representations of $G$. We study the asymptotic concentration of the eigenfunctions in a given isotypical component.
@@ 1 Introduction
As proved and discussed in [GStz-1991], [GStz-1992], [Lem], [LS], [SZ02] (see also [PW-1989] and [PW-1991]), the choice of a $\mathcal{C}^{\varpi}$ Riemannian metric $\kappa$ on $M$ singles out a unique $\mathcal{C}^{\varpi}$ exhaustion function on $\tilde{M}$ (perhaps after passing to a smaller tubular neighbourhood of $M$ in $\tilde{M}$), with the following properties:
@@
Grauert tubes have been studied extensively in recent years, both for their intrinsic geometric interest and in relation to the a fundamental holomorphic extension property of the eigenfunctions of the non-negative Laplacian $\Delta$ of $(M,\kappa)$, which was first discovered by Bouted de Monvel [BdM-1978]; this foundational result was also proved and analyzed in [Leb], [S-2014], [S-2015],[Z-2020] (see furthermore the discussion in [GStz-1992]). Grauert tube techniques have also been applied to the study of nodal sets (see [CT-2016], [CT-2018], [TZ-2009], [TZ-2021]).
Let us recall Boutet de Monvel’s foundational result, as formulated in [Leb]. Denote by
@@
    $\left(e^{-\tau\,\mu_{j}}\,\langle\mu_{j}\rangle^{(d-1)/4}\,\tilde{\varphi}_{j,k}^{\tau}\right)_{j\geq 1,\,k=1,\ldots,n_{j}}$
@@
    a privileged collection of CR functions (elements of $H(X^{\tau})$), given by the restrictions $\tilde{\varphi}^{\tau}$ to $X^{\tau}$ of the holomorphic extensions $\tilde{\varphi}$ of the eigenfunctions $\varphi$ of $\Delta$ (if $\tau\in(0,\tau_{1})$);
@@
In fact, the principal symbol of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ (i.e., the restriction to $\Sigma^{\tau}$ in (6) of the principal symbol of $D^{\tau}_{\sqrt{\rho}}$) is strictly positive; hence the spectrum of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$ is discrete and accumulates only at $\infty$ [BdM-G].
@@
Let us notice the following regarding the two conditions in Assumption 4.
  1. 1.
    The second requirement is equivalent to $0\in\mathfrak{g}^{\vee}$ being a regular value of $\left.\Phi\right|_{\tilde{M}^{\tau_{0}}\setminus M}$ (see e.g. [$M$], [GStb]); therefore, it implies that $Z\setminus M$ and $Z^{\tau}$ are submanifolds of $\tilde{M}$ and $X^{\tau}$, respectively. Our scaling asymptotics rely on a direct sum decomposition of $TX^{\tau}$ along $Z^{\tau}$ that depends on this smoothness assumption (see §1.1.2).
  2. 2.
    $\tilde{\mu}^{\tau}$ is a fortiori locally free on $Z^{\tau}$ if $\mu$ itself is locally free, e.g., if $M$ is a principal $G$-bundle; for more examples, see §2.3.
  3. 3.
    Under Assumption 4, $d>d_{G}$ (see Corollary 55 below); for instance, the case of a compact Lie group acting on itself by, say, left translations is not covered by the present analysis. This is because $Z^{\tau}=\emptyset$ in this case (we shall consider this specific situation in a separate paper).
Let $\Gamma:\mathbb{R}\times\tilde{M}\rightarrow\tilde{M}$ denote the flow of $\upsilon_{\sqrt{\rho}}$ ($\Gamma$ will be referred to - with some abuse of language - as the ‘homogeneous geodesic flow’,   since it is intertwined with the latter by the imaginary time exponential map - see [GStz-1992], [LS], [GLS]). Its restriction to $X^{\tau}$ is the flow of $\upsilon_{\sqrt{\rho}}^{\tau}$, and will be denoted $\Gamma^{\tau}:\mathbb{R}\times X^{\tau}\rightarrow X^{\tau}$. Since $\tilde{\mu}^{\tau}$ and $\Gamma^{\tau}$ commute, there is a product action of $G\times\mathbb{R}$ on $X^{\tau}$; our first result is that (16) and (17) asymptotically concentrate near certain compact loci $\mathfrak{X}^{\tau}_{\chi}\subset X^{\tau}\times X^{\tau}$ determined by $\chi$, the moment map, and the orbits of the latter action.
@@
We are thus led to consider the asymptotics of $P^{\tau}_{\chi,\nu,\lambda}(x,y)$ and $\Pi^{\tau}_{\chi,\nu,\lambda}(x,y)$ when $(x,y)$ ranges in a shrinking neighourhood of a fixed pair $(x_{1},x_{2})\in\mathfrak{X}^{\tau}_{\chi}$.
1.1 Scaling asymptotics
The previous estimates motivate the expectation that near $\mathfrak{X}^{\tau}_{\chi}$ the two kernels satisfy scaling asymptotics on a scale of $O(\lambda^{-1/2})$ exhibing an exponential decay along directions transverse to $\mathfrak{X}^{\tau}_{\chi}$. The aim of the following statements is to substantiate this expectation.
Before giving precise statements, it is in order to premise a few general remarks. While the general approach is heuristically inspired by the line bundle setting, the arguments are significantly more involved in the present context, the primary reason being that the geodesic flow is generally not holomorphic. As the reader will appreciate, Zelditch’ method of dynamical Toeplitz operators plays an essential conceptual and technical role in dealing with this difficulty. Another key ingredient is the use of suitable sets of local coordinates (whose construction goes back to Folland and Stein in [FS1] and [FS2], and which were first used in this context by Chang and Rabinowitz), specifically adapted to the local CR geometry of $X^{\tau}$.
More precisely, as in [P-2024] the asymptotics in point will be formulated in suitable systems of local coordinates at points $x\in X^{\tau}$, called normal Heisenberg local coordinates (in the following, NHLC’s), which are a slight specialization of the Heisenberg local coordinates in [CR1] and [CR2] (see §3.3 of [P-2024]).
@@ 1.1 Scaling asymptotics
The non-holomorphicity, or equivalently the non-unitarity, of the geodesic flow is encapsulated in the appearance of a (generally) non-unitary matrix $B$ in the local description of $\Gamma^{\tau}$ (see (25) and §2.6.2), and in the ensuing oscillatory and Gaussian integrals computing the asymptotics.
In order to obtain tractable computations and extract more intelligible geometric information, we have in some cases restricted the type of rescaled directions. Specifically, in the near-graph action-free setting of [CR2] (Theorem 15 below) we consider arbitrary rescaled displacements; similarly, in the near-diagonal equivariant case (Theorem 28) we allow arbitrary displacements away from $G$-orbits. However, in the general equivariant case we consider more specific choices of directions (see below).
1.1.1 Scaling asymptotics in the action-free case
@@
The leading order term describing the exponential decay in the scaling asymptotics of $P^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ and $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ for $\lambda\rightarrow+\infty$ depends on a real-quadratic complex-valued form on the vector subspace $\mathcal{H}_{x_{1}}^{\tau}\times\mathcal{H}_{x_{2}}^{\tau}\subseteq T_{x_{1}}X^{\tau}\times T_{x_{2}}X^{\tau}$. To describe the latter form, we need a brief digression.
@@
$:=$ $\frac{1}{2}\,\left(Z_{1}^{\dagger}\,\overline{Q}\,P^{-1}\,Z_{1}+2\,\overline{Z}_{2}^{\dagger}\,P^{-1}\,Z_{1}-\overline{Z}_{2}^{\dagger}\,P^{-1}\,Q\,\overline{Z}_{2}-\|Z_{1}\|^{2}-\|Z_{2}\|^{2}\right).$
@@
If $(V,h^{\prime})$ is a Hermitian complex $k$-dimensional vector space, so that $h=g_{h}-\imath\,\omega_{h}$, where $g_{h}:=\Re(h)$ and $\omega_{h}:=-\Im(h)$ are, respectively, an Euclidean product and a symplectic bilinear form on $V$. We define $\psi_{2}^{\omega_{h}}:V\times V\rightarrow\mathbb{C}$ by
@@
Let us identify the unitary group U(d1)U(d-1) as the maximal compact subgroup $\hat{U}(d-1):=\mathrm{Sp}(2d-2)\cap O(2d-2)$ of $\mathrm{Sp}(2d-2)$ in the standard manner. Then $A\in\hat{U}(d-1)\leqslant\mathrm{Sp}(2d-2)$ if and only if in (22) we have $Q=0$ and $P\in U(d-1)\leqslant\mathrm{GL}_{\mathbb{C}}(d-1)$. Hence if $A\in\hat{U}(d-1)$ then
$\Psi_{A}(\mathbf{v}_{1},\mathbf{v}_{2})$ $=$ $\overline{Z}_{2}^{\dagger}\,P^{-1}\,Z_{1}-\frac{1}{2}\,\|Z_{1}\|^{2}-\frac{1}{2}\,\|Z_{2}\|^{2}$
$=$ $h_{0}(Z_{1},P\,Z_{2})-\frac{1}{2}\,\|Z_{1}\|^{2}-\frac{1}{2}\,\|Z_{2}\|^{2}$
@@
When $x_{1}=x_{2}$ and $\chi\in\mathcal{C}^{\infty}_{c}\big{(}(-\epsilon,\epsilon)\big{)}$ for suitably small $\epsilon>0$ then $M=I_{2d-2}$, $t_{1}=0$ and $\theta^{\tau}(x_{1},x_{2})=0$ Thus we recover the near-diagonal scaling asymptotic in [P-2024]. Furthermore, we recover the near-graph scaling asymptotics of [CR2], with an explicit determination of the leading factor, by fixing $(\theta_{j},\mathbf{v}_{j})$ and rescaling according to Heisenberg type (that is, in the form $(\theta_{j}/\lambda,\mathbf{v}_{j}/\sqrt{\lambda})$). The analogue of Theorem 15 for $P^{\tau}_{\chi,\nu,\lambda}$ is discussed in Section 1.1.4.
1.1.2 A decomposition of $T_{x}X^{\tau}$ in the general equivariant case
In order to state our results in the general equivariant setting, we need to introduce a finer decomposition of $T_{x}X^{\tau}$, valid at $x\in Z^{\tau}$, than the one dictated by $\alpha^{\tau}$ in (8); this decomposition depends on $\mu^{\tau}$ and $\Phi$, and is the analogue of the decomposition appearing in the equivariant asymptotics in the line bundle setting (see [MZ] and [P-2008]).
@@ 1.1.2 A decomposition of $T_{x}X^{\tau}$ in the general equivariant case
Given a choice of NHLC’s centered for $X^{\tau}$ at $x$, we obtain a direct sum decomposition of $\mathbb{R}^{2\,d-1}=\mathbb{R}\oplus\mathbb{C}^{d-1}\cong T_{x}X^{\tau}$, dictated by $\alpha^{\tau}$, where $\mathbb{C}^{d-1}\cong\mathcal{H}^{\tau}_{x}$. The finer decomposition of $\mathcal{H}^{\tau}_{x}$ in Definition 18, dictated by $\tilde{\mu}^{\tau}$, determines a corresponding decomposition of $\mathbb{C}^{d-1}$, which - emphasizing dimensions - we shall write in the form
@@ 1.1.3 Scaling asymptotics for $\Pi^{\tau}_{\chi,\nu,\lambda}$
Under Assumption 4, the effective volume at $x\in Z^{\tau}$, denoted $V_{eff}(x)$, is the volume of the $G$-orbit through $x$, $x^{G}\subset X^{\tau}$ for the induced Riemannian density (with respect to $\tilde{\kappa} -$ see also Remark 14).
@@
$A_{\chi}(x_{1},x_{2})$ $:=$ $\int_{\mathbb{R}^{d_{G}}_{t}}\,\mathrm{d}\mathbf{u}^{t}\,\int_{\mathbb{R}^{d_{G}}_{v}}\,\mathrm{d}\mathbf{u}^{v}\,\int_{\mathbb{R}^{2d-2-2d_{G}}_{h}}\,\mathrm{d}\mathbf{u}^{h}$ (26)
@@
$\mathcal{F}_{\chi}(x_{1},x_{2})$ $:=$ $\chi(t_{1})\cdot\frac{\left|\det(P)\right|}{r\cdot V_{eff}(x_{1})}\cdot\frac{A_{\chi}(x_{1},x_{2})}{\pi^{d-1}},$
$\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$ $:=$ $\dim(\nu)\cdot\overline{\Xi_{\nu}\left(g_{l}\right)},$
@@
    if $B$ is orthogonal one also has $|\det(P)|=1$ (since in this case $P$ is a unitary complex matrix), hence $\mathcal{F}_{\chi}(x_{1},x_{2})=\chi(t_{1})/\big{(}r\cdot V_{eff}(x_{1})\big{)}$.
In view of Theorem 7, we expect an exponential decay of $\Pi^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$ (and $P^{\tau}_{\chi,\nu,\lambda}(x_{1,\lambda},x_{2,\lambda})$) along normal displacements to $Z^{\tau}$, that is, when $\mathbf{v}_{j}=\mathbf{v}_{j}^{t}\in T_{x_{j}}^{t}X^{\tau}$, and we wish to determine the leading exponent governing. Accordingly, in the general equivariant case we restrict first to dispacements along directions in $T_{x_{j}}^{t}X^{\tau}$ (Theorem 25, case 1); in this case, the leading exponent is determined explicitly. In the same setting, an expansion is also obtained, although with a less explicit determination of the exponent, for more general displacements, provided we take $\mathbf{v}_{2}^{h}=0$ (Theorem 25, case 2). As discussed at the beginning of §1.1, this unequal treatment of $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ is aimed at making the computations more tractable, while obtaining managable results that are sufficient for the applications.
@@
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{k\geq 1}\,\lambda^{-k/2}\,F_{k,l,\nu}\left(x_{1},x_{2};\theta_{1},\mathbf{v}_{1}^{t},\theta_{2},\mathbf{v}_{2}^{t}\right)\right];$
@@
    $\mathbf{v}_{1}=\mathbf{v}_{1}^{t}+\mathbf{v}_{1}^{h}\in T^{t}_{x_{1}}X^{\tau}\oplus T^{h}_{x_{1}}X^{\tau},\quad\|\mathbf{v}_{1}\|\leq C\,\lambda^{\epsilon^{\prime}}$
@@
    $\cdot\left[e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}+\sum_{k\geq 1}\,\lambda^{-k/2}\,\tilde{F}_{k,l,\nu}\left(x_{1},x_{2};\theta_{1},\theta_{2},\mathbf{V}\right)\right];$
@@
The unitary factor $e^{\imath\,\theta^{\tau}_{t}(x)}$ is a smooth function on $\mathbb{R}\times X^{\tau}$ and is related to the description (by Zelditch) of the unitary Toeplitz operators $e^{\imath\,t\,\mathfrak{D}^{\tau}_{\sqrt{\rho}}}$ as ‘dynamical Toeplitz operators’ (recall (10)); it is therefore an intrinsic invariant, given a priori, of the CR structure of $X^{\tau}$. Similar considerations apply to the unitary factor $e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}$ appearing in the asymptotics of $P^{\tau}_{\chi,\nu,\lambda}$ of §1.1.4, with $e^{\imath\,t\,\mathfrak{D}^{\tau}_{\sqrt{\rho}}}$ replaced by $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ in (6) below - see (162) and (164) below.